OpenStudy (anonymous):

An athlete can burn 10 calories per minute while running and 4 calories per minute while walking. If the athlete burns 450 calories in 60 minutes, how long is spent on each activity?

OpenStudy (hoblos):

x=running time y=walking time x+y=60 10x+4y=450 can you solve the system??

OpenStudy (anonymous):

I thought t needs one variable?

OpenStudy (hoblos):

i'll solve it for you first multiply the first equation by 4 4x+4y=240 10x+4y=450 subtract -6x = -210 x=35 min y=60-35 =25 min